Please help me solve this problem: 4-6y 4=4y 16-6y.
Found 2 solutions by fractalier, algebra hello:
Answer by fractalier(6550)

Okay from
4-6y+4=4y+16-6y
we gather like terms and get
-6y + 8 = -2y + 16
Now subtract 8 from both sides
-6y = -2y + 8
And now add 2y...
-4y = 8
Now divide by -4...we get
y = -2

Answer by algebra hello(55)

on writting the variable terms on left side and constant terms on other side, we get
-6y-4y 6y=16-4-4
-4y=8
y=-8/4
y=-2